One number is 8 less than another. If 5 times the smaller number minus 2 times the larger number is 14, find the two numbers. I'm confused, Please help...
Answer by checkley75(3666)

X=Y-8
5X-2Y=14 NOW SUBSTITUTE (Y-8) FOR X & SOLVE FOR Y
5(Y-8)-2Y=14
5Y-40-2Y=14
3Y=54
Y=54/3
Y=18 THE LARGER NUMBER.
X=18-8
X=10 FOR THE SMALLER NUMBER.
PROOF
5*10-2*18=14
50-36=14
14=14